# Learning-based Multi-View Stereo: A Survey

### v1

@@ V-E5 Real-time and Memory Efficient Models
Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving. Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability. Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [31, 35]. In addition, the efficiency in both run-time and memory will drop when image resolution increases. For offline MVS methods, as discussed in Sec. V-B2, some recent methods try to improve efficiency with carefully simplified network architectures. However, this usually limits the performance when compared with state-of-the-art methods. There are many other directions for further improvement. For example, using model compression techniques and knowledge distillation are potential directions to not only keep the good performance and also improve efficiency.
V-E6 Prior Assistance
@@
MVS mainly relies on measurements of local photometric consistency to find the optimal matching across reference and source images. Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made environments, e.g., ETH3D [23]. Therefore, leveraging prior information to guide the MVS algorithm in these challenging regions is a promising research direction. We elaborate several typical examples of prior assistance as follows.
@@
  • Surface Normal: It is a non-local representation of the geometry compared with the depth map. To enforce the constraints of normal maps in depth estimation, Kusupati et al. [206] integrate the multi-view normal estimation network into the MVS pipeline. Long et al. [207] introduce a Combined Normal Map, estimated with PlaneCNN [208] to enforce the normal consistency on the depth map. Liu et al. [209] perform graph-based depth map optimization as post-processing. Since normal monocular normal predictions [210] provide high-quality priors for the full scene, [154, 155] use monocular normal to improve the optimization of neural implicit surfaces.
@@
  • Semantic Segmentation: Intuitively, points assigned with the same semantic labels may be more likely to lie on the same 3D plane. Some attempts [214, 215] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality. Manhattan-SDF [216] adopts Manhattan world priors to handle low-textured planar regions corresponding to walls and floors with 2D segmentation.
VI Conclusion
In this survey, we have provided a comprehensive review of the learning-based MVS methods until 2023, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability. We explain key aspects such as datasets utilized, general working pipelines, and algorithmic intricacies. Furthermore, we have summarized and compared the quantitative performance of these methods across popular benchmarks, providing valuable insights into their efficacy and applicability. Finally, our discourse extends to an in-depth exploration of potential research directions for the future of MVS, highlighting avenues for further investigation and innovation in the field.

### v2

@@ 6.5.5 Real-time and Memory Efficient Models
Though reconstruction accuracy is considered the most important factor in MVS, it is also critical to have models that can run in near real-time and with low memory, e.g., AR/VR and autonomous driving. Comparatively, depth map-based methods are most suitable for achieving efficiency because of their conciseness, flexibility, and scalability. Though online MVS methods can achieve real-time estimation with images of low resolutions, they may have issues with memory consumption since many large backbones are usually used for feature extraction [30, 34]. In addition, the efficiency in both run-time and memory will drop when image resolution increases. For offline MVS methods, as discussed in Sec. 6.2.2, some recent methods try to improve efficiency with carefully simplified network architectures. However, this usually limits the performance when compared with state-of-the-art methods. There are many other directions for further improvement. For example, using model compression techniques and knowledge distillation are potential directions to not only keep the good performance and also improve efficiency.
6.5.6 Prior Assistance
@@
MVS mainly relies on assessing local photometric consistency to find the optimal matching across reference and source images. Accordingly, it usually encounters difficulties when estimating the geometry for regions where the photometric measurement becomes unreliable or invalid, e.g., textureless areas and non-Lambertian surfaces, which are common in human-made scenes [23]. Therefore, using prior information to guide the MVS algorithm in these challenging regions is a promising research direction. We elaborate 3 typical examples of prior assistance as follows.
@@
Surface Normal: As a non-local representation of the geometry compared with the depth map, surface normal has been proven effective in improving reconstruction performance in recent works. To enforce the constraints of normal maps in depth estimation, Kusupati et al. [207] integrate a multi-view normal estimation network into the MVS pipeline. Long et al. [208] introduce a Combined Normal Map, estimated with PlaneCNN [209] to enforce the normal consistency on the depth map. Liu et al. [210] apply graph-based depth map optimization as post-processing. Since monocular normal network [211] is trained on large-scale dataset and can provide high-quality priors, [146, 187] use monocular normal priors to improve the surface accuracy.
@@
Semantic Segmentation: Intuitively, points assigned with the same semantic labels may be more likely to lie on the same 3D plane. Some attempts [215, 216] employ a rule-based protocol to utilize semantic segmentation for better matching and depth quality. Manhattan-SDF [217] relies on 2D segmentation results to enforce Manhattan world priors to handle low-textured planar regions. Similarly, Shvets et al. [218] also employ SAM (Segment Anything Model) encoder to generate semantic features to enhance MVS.
7 Conclusion
In this survey, we have provided a comprehensive review of the learning-based MVS methods, which are categorized into: depth map-based, voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods. Particularly, we have devoted significant attention to depth map-based methods because of their conciseness, flexibility and scalability. We explain key aspects such as datasets utilized, general working pipelines, and algorithmic intricacies. Furthermore, we have summarized and compared the quantitative performance of these methods across popular benchmarks, providing valuable insights into their efficacy and applicability. Finally, our discourse extends to an in-depth exploration of potential research directions for the future of MVS, highlighting avenues for further investigation and innovation in the field.